Shift Summary for OWL, June 10, 1999

Alexander Ledovskoy (ledovsko@fnpxsr)
Thu, 10 Jun 1999 07:02:27 -0500

Shift Summary for OWL shift, 10-June-99

On Shift: D. Jensen
S. Ledovskoy

Also: Miguel, Elizabeth, Rick K.

Hours of beam delivered at all intensities:
Hours of beam delivered at nominal intensity:
Hours of E832 data (special runs not included):
Hours of special runs (laser scans, muon runs...):

Accounting of lost time:

========================================================================
Runs ended this shift:

13473 407K events BEAM_MU_4E12
13474 283K events BEAM_MU_4E12
13475 227K events BEAM_MU_4E12
13476 217K events BEAM_MU_4E12
13477 109K events BEAM_MU_4E12
13478 137K events BEAM_MU_4E12
13479 334K events BEAM_MU_4E12
13480 198K events BEAM_MU_4E12
13481 99K events BEAM_MU_4E12
13482 170K events BEAM_MU_4E12
13483 233K events BEAM_MU_4E12
13484 182K events BEAM_MU_4E12

Tapes brought to FCC this shift:

========================================================================
Accesses during shift:

========================================================================
Overview of shift:

Muon runs all shift long. There is no hadrons due to NM2S3 power supply
failure.

Miguel studies DC plateau
Elizabeth works on MA and REG HV
Rick K works on filter code and trigger configrations

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
13473 00:00 BEAM_MU_4E12 is running 10pm yesterday
109 00:28 Stop the run

Beam Control called and asked about our plans.
We informed them that we will ask 2E12 beam ,
11 batches later this shift. They expressed
concern that they will not be able to do that.
It is more likely that intensity will be lower
at 3 batches only

13474 00:40 BEAM_MU_4E12 started

We've been taking muon runs for a while. It is
possible that muon data will fill up the disks and
the button "DISK" turns red in ONLINE SPLIT window.
If this happens, page Sydney (722-6849)

00:50 Stop the run

13475 01:00 BEAM_MU_4E12 started

01:05 Main Control called. There is an indication of
water leak in NM2S3. Beam is off. They are investi-
gating.

01:15 Gas Checklist is done.

01:27 It is verified, there is an internal water leak
in power supply of NM2S3. Repairs will start
in the morning. Before that we will have muons.
There will be no hadrons.

01:50 The run stopped

13476 01:55 BEAM_MU_4E12 started

"CACHER" button turns red in ONLINE SPLIT window.
It goes green few minutes later. This is a false
alarm according to Sydney (who was paged anyways)

02:03 The run is stopped

13477 02:05 BEAM_MU_4E12 started

13478 02:20 BEAM_MU_4E12 started

13479 02:35 BEAM_MU_4E12 started

02:40 Main Control called and told us that they've been
delivering beam in 11 batches for last 0.5 hours.

13480 02:45 BEAM_MU_4E12 started

13481 03:15 BEAM_MU_4E12 started

13482 03:27 BEAM_MU_4E12 started

13483 03:40 BEAM_MU_4E12 started

13484 03:50 BEAM_MU_4E12 started

Miguel finished DC studies

13487 04:50 BEAM_MU_4E12 started.
This is going to be long muon run
(goal is to get 2M events)

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

1. The alignment crew for DC will be here at 8:00am
2. Get in touch with Brian Lavoy as soon as alignment crew is here.
3. get 4413 CAMAC discriminator from PREP (call x3447 first)